### v2

# Quantitative Evaluation of Multiple Instance Learning Reliability For WSIs Classification

Abstract
[…]
keywords:
Machine learning, Reliability, Whole Slide Image, Multiple Instance Learning, MEAN-POOL-INS.
1 Introduction
Multiple Instance Learning (MIL) is a widely used method for weakly-supervised classification tasks in computational pathology, primarily due to the difficulty and labor-intensive nature of obtaining pixel-wise annotations from histopathology data. In the MIL framework, groups of instances, or "bags," are created, with the goal of predicting the label for each bag. The key assumption in MIL is that positive bags contain at least one key instance, while negative bags do not. Recent advancements in deep learning have spurred the development of novel MIL models tailored to computational pathology, enhancing performance through key instance detection [Ilse et al.(2018)Ilse, Tomczak, and Welling], feature space refinement [Lu et al.(2021)Lu, Williamson, Chen, Chen, Barbieri, and Mahmood], and overfitting mitigation [Zhang et al.(2022)Zhang, Meng, Zhao, Qiao, Yang, Coupland, and Zheng], [Zhang et al.(2023)Zhang, Li, Sun, Zheng, Zhu, and Yang].
Reliability refers to the ability of a system or model to perform as expected under standard and well-defined conditions. The reliability of MIL models is essential for gaining the trust of physicians, ensuring successful deployment in clinical settings, and adhering to established healthcare standards. Reliable models focus on biologically pertinent features, grounded in established scientific knowledge, when making predictions. For instance, in the context of computational pathology, an unreliable MIL model might erroneously focus on artifacts or irrelevant features, leading to false positives or negatives in critical tasks such as tumor detection or subtyping. Such errors could result in misdiagnoses, delays in treatment, or unnecessary interventions, underscoring the need for robust reliability evaluations. Achieving reliability in MIL models is particularly challenging due to the inherent variability in histopathology data, including differences in staining, scanning resolutions, and patient-specific biological heterogeneity. Furthermore, the weakly-supervised nature of MIL exacerbates these challenges by relying on coarse slide-level labels without precise spatial annotations. Addressing these limitations requires not only model improvements but also careful evaluation frameworks that ensure predictions align with domain knowledge and clinical relevance. By fostering reliable MIL models, computational pathology can advance toward broader acceptance and integration into clinical workflows.
However, machine learning models are rarely evaluated in terms of reliability; instead, performance metrics such as accuracy and F1 score are typically reported. Reliable models are easier to deploy in real-world applications and tend to be more generalizable. Therefore, gaining insights into models’ reliability is critical. For instance, Pendyala et al. [Pendyala and Kim(2024)] assessed the reliability of various models using a mental health dataset and discovered that these models often focused on less relevant features, despite achieving high performance, thereby highlighting the limitations of traditional metrics. Furthermore, Araujo et al. [Araujo Santos et al.(2023)Araujo Santos, Cardoso, and Alves] propose an alternative validation method for models and support the assertion that performance metrics are not indicative of the reliability of machine learning networks. In the domain of medical imaging, the trustworthiness of saliency maps has been quantitatively assessed on two public radiology datasets using the area under the precision-recall curve (AUPRC) and the structural similarity index measure (SSIM) as evaluation metrics [Arun et al.(2021)Arun, Gaw, Singh, Chang, Aggarwal, Chen, Hoebel, Gupta, Patel, Gidwani, et al.]. The study revealed that none of the saliency maps satisfied all the tested criteria, highlighting limitations in their reliability.
Interpretability and reliability in the context of machine learning (ML) are related but distinct concepts that address different aspects of model performance and trustworthiness. We aim to understand and explain the way models work. To address questions concerning the decision-making processes and the underlying mechanics, numerous interpretability methods have been developed [Salahuddin et al.(2022)Salahuddin, Woodruff, Chatterjee, and Lambin]. Interpretability and explainability methods have proven useful in identifying both data and model deficiencies [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash]. For MIL models, attention scores assigned to each patch are frequently used to generate heatmaps, which are key tools for assessing interpretability. However, interpretability in MIL models has traditionally been evaluated qualitatively by presenting specific slides and their corresponding heatmaps. The Camelyon16 dataset [Bejnordi et al.(2017)Bejnordi, Veta, Van Diest, Van Ginneken, Karssemeijer, Litjens, Van Der Laak, Hermsen, Manson, Balkenhol, et al.], which includes pixel-wise ground truth annotations, is the most common resource for such evaluations. While qualitative assessments provide valuable insights, they are not feasible across large test sets due to computational and time constraints. Moreover, analyzing heatmaps requires specialized histopathology knowledge, which may not be readily available to machine learning researchers. Finally, a more robust and comprehensive assessment should involve evaluation across multiple datasets.
Reliability is concerned with ensuring the model consistently produces correct and trustworthy predictions in a wide range of circumstances, independent of transparency of underlying process by which predictions are made. Models focusing on related regions of interest (ROIs) are more reliable, as they process the most relevant parts of the data. However, more interpretable models are not necessarily more reliable. For instance, a model might offer more interpretability through class-specific information or excitatory-inhibitory heatmaps [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash], yet fail to concentrate on relevant ROIs. Additionally, while interpretable models may provide explanations for predictions across entire classes, this does not necessarily align with reliability. For example, in binary tumor versus normal classification, since normal samples lack clear ROIs, predictions for these cases are not considered in reliability assessments.
In this paper, we address the above-mentioned challenges by quantitatively evaluating the reliability of MIL models for both binary tumor versus normal classification and multi-class tissue subtyping problems using three metrics. Additionally, we integrate the recently proposed additive method [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash], designed to enhance MIL model interpretability, and evaluate if this also impacts reliability. To provide a more comprehensive assessment, we extend our experiments beyond the Camelyon16 dataset by including two additional datasets: CATCH [Wilm et al.(2022)Wilm, Fragoso, Marzahl, Qiu, Puget, Diehl, Bertram, Klopfleisch, Maier, Breininger, et al.] and TCGA BRCA. Our study aims to advance the development of more reliable MIL models for computational pathology.
2 Method
2.1 Problem formulation
For our analysis, we refer to the input Whole Slide Image (WSI) as $I$ and predict its slide-level label $Y$. Following established methodologies, we first tessellate the slide into $N$ small patches and then extract features from these patches using a pre-trained model. In the final step, the slide label is predicted using an aggregation module, which is typically attention-based, providing attention scores for each patch. For clarification, the proposed framework is illustrated in Figure LABEL:fig:anno.
2.2 Metrics
We propose using three metrics to quantitatively evaluate the reliability of MIL models for WSI classification. Specifically, since our aim is to verify whether MIL models focus on ROIs, we evaluate the alignment between patch scores and annotations from three distinct perspectives:
  • Mutual Information (MI): MI is a measure of statistical dependence. It has been widely used for feature selection and transformation in machine learning [Vergara and Estévez(2014)]. MI quantifies how well values in image $X$ at a given location can predict corresponding values in $Y$. For our purpose, we use MI to evaluate the correspondence of patch scores with their corresponding ground truth.
  • Spearman’s Correlation $(r_{s})$: Spearman’s correlation measures the strength and direction of a non-linear association between two variables. We apply this metric to assess the non-linear correlation between predicted patch scores and ground truth annotations.
  • Area Under The Precision-Recall Curve $(AUPRC)$: AUPRC is particularly useful for evaluating models on highly imbalanced datasets. Given that often our patch annotations are imbalanced, we use AUPRC to measure the localization performance effectively.
3 Experiment and Results
3.1 Dataset
Three public WSI datasets were used in the experiments. The Breast Invasive Carcinoma dataset (TCGA BRCA) from The Cancer Genome Atlas (TCGA) project includes 1,038 H&E-stained slides representing two types of breast cancer: Invasive Ductal Carcinoma (IDC) and Invasive Lobular Carcinoma (ILC), with slide-level labels. The CAMELYON16 dataset contains 399 WSIs with breast cancer lymph node metastasis annotations, divided into 270 slides for training and validation, and 129 for testing. Finally, the CATCH dataset provides 350 meticulously annotated slides from seven different canine cutaneous tumors.
Unlike CAMELYON16 and CATCH, the TCGA BRCA dataset lacks ground truth region-level annotations. To address this, tumor regions in the test slides were manually annotated with the assistance of Dr. Christof A. Bertram, a pathologist.
3.2 Implementation details
Throughout the experiments, we employ the Adam optimizer with a learning rate of 1e-4 and set the maximum number of epochs to 50. The weight decay is tuned based on the validation loss. All experiments are repeated with five different random seeds, and we report the mean AUC and F1 scores as classification metrics, along with the mean MI, $r_{s}$, and AUPRC values as reliability metrics. Additionally, we evaluate model size and Floating-Point Operations Per Second (FLOPS) as computational metrics. For dataset preprocessing, we follow the procedures outlined in [Lu et al.(2021)Lu, Williamson, Chen, Chen, Barbieri, and Mahmood], which involve extracting features from non-overlapping 256×256 patches at 20× magnification using a ResNet50 pretrained on ImageNet. These extracted features are fed into the model with an input dimensionality of $input\_dim=1024$, which is subsequently reduced to $D=512$.
We evaluate several state-of-the-art models, including ABMIL [Ilse et al.(2018)Ilse, Tomczak, and Welling], CLAM [Lu et al.(2021)Lu, Williamson, Chen, Chen, Barbieri, and Mahmood], DTFD [Zhang et al.(2022)Zhang, Meng, Zhao, Qiao, Yang, Coupland, and Zheng], ACMIL [Zhang et al.(2023)Zhang, Li, Sun, Zheng, Zhu, and Yang], MADMIL [Keshvarikhojasteh et al.(2024)Keshvarikhojasteh, Pluim, and Veta], mean pooling (MEAN-POOL), max pooling (MAX-POOL), max pooling instance (MAX-POOL-INS), and mean pooling instance (MEAN-POOL-INS). For the MEAN-POOL and MAX-POOL models, we calculate the slide representation after feature compression using either the mean or max operation, and then predict the slide label. In contrast, for the MAX-POOL-INS and MEAN-POOL-INS models, we first predict patch probabilities, which are then aggregated to generate slide-level probabilities. Patch probabilities are also used to compute reliability metrics for MEAN-POOL-INS. In the case of MAX-POOL-INS, we use both the selected patch index, similar to MAX-POOL, and the patch probabilities.
For multi-head models such as MADMIL and ACMIL, average attention scores or patch probabilities are employed to calculate reliability metrics. We use a shorthand notation to indicate the number of heads, e.g., MADMIL/4 refers to the MADMIL model with four heads. To ensure a thorough analysis, we also investigate additive versions of these models, in accordance with the approach proposed by [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash] , which reformulates the final predictor as an additive function of individual instances. For additive models, either attention scores or patch probabilities are used to assess reliability.
To compute the reliability metrics, it is essential to obtain both patch labels and the corresponding attention scores or probabilities. Following tessellation, patch labels are assigned based on the region-wise ground truth annotations. Specifically, for each patch, if it falls within a annotated region , it is assigned a label of 1, indicating the presence of the target feature. Conversely, if the patch does not intersect with the annotated region, it is assigned a label of 0, denoting the absence of the target feature.
[…]
fig:anno Refer to caption
Figure 1: The overall framework for evaluating the reliability of MIL models follows a three-step process. First, a MIL model is trained on a weakly-supervised task for predicting slide-level labels. Next, the trained model is applied to predict scores for individual image patches. Finally, the reliability value is computed based on the predicted patch scores and their corresponding annotations, where tumor patches are highlighted in green and normal patches in orange in the annotation visualization.
Table 1: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CAMELYON16 dataset. FLOPs are calculated using 120 instances per bag.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.1166 ± 0.0048 0.2882 ± 0.0044 0.5850 ± 0.0101 0.847 ± 0.011 0.844 ± 0.007 94.4 M 788.7 K CLAM 0.1176 ± 0.0089 0.2925 ± 0.0173 0.5889 ± 0.0220 0.860 ± 0.005 0.850 ± 0.007 94.4 M 790.7 K DTFD 0.1275 ± 0.0071 0.2939 ± 0.0106 0.6057 ± 0.0142 0.893 ± 0.011 0.847 ± 0.017 125.8 M 1052.6 K ACMIL/2 0.1612 ± 0.0198 0.3352 ± 0.0168 0.5856 ± 0.0486 0.824 ± 0.028 0.810 ± 0.014 94.4 M 791.0 K MADMIL/2 0.1330 ± 0.0045 0.3057 ± 0.0043 0.6080 ± 0.0037 0.862 ± 0.009 0.847 ± 0.013 78.6 M 657.6 K MAX-POOL 0.0018 ± 0.0001 -$0$ 0.0014 0.2253 ± 0.0097 0.798 ± 0.006 0.793 ± 0.012 62.9 M 525.8 K MEAN-POOL [^1] The reliability metrics are not computed for this model, as it does not provide patch scores. - - - 0.581 ± 0.022 0.589 ± 0.021 62.9 M 525.8 K
[…]
fig:cam_max Refer to caption        Refer to caption     Refer to caption
Figure 2: The test-30 slide with ground truth annotation (green) along with corresponding heatmap generated by MAX-POOL.
Table 2: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CAMELYON16 using additive models.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.1336 ± 0.0031 0.3179 ± 0.0081 0.6152 ± 0.0107 0.861 ± 0.010 0.827 ± 0.015 94.5 M 788.7 K ABMIL-ADD-PATCH 0.0380 ± 0.0025 0.2377 ± 0.0112 0.4848 ± 0.0070 CLAM-ADD-ATT 0.1153 ± 0.0091 0.2924 ± 0.0089 0.5968 ± 0.0170 0.865 ± 0.006 0.826 ± 0.019 94.5 M 790.7 K CLAM-ADD-PATCH 0.0325 ± 0.0042 0.2394 ± 0.0499 0.4826 ± 0.0165 DTFD-PATCH  [^2] We use the provided patch probabilities based on GRAD-CAM. 0.0368 ± 0.0030 0.2895 ± 0.0197 0.5308 ± 0.0124 0.893 ± 0.011 0.847 ± 0.017 125.8 M 1052.6 K ACMIL/2-ADD-ATT 0.1739 ± 0.0055 0.3463 ± 0.0117 0.6201 ± 0.0220 0.787 ± 0.024 0.810 ± 0.011 94.5 M 791.0 K ACMIL/2-ADD-PATCH 0.0263 ± 0.0033 -0.0346 ± 0.0120 0.2873 ± 0.0088 MAX-POOL-INS 0.0004 ± 0.0002 0.0473 ± 0.0255 0.1986 ± 0.0228 0.766 ± 0.135 0.763 ± 0.127 63.0 M 525.8 K MAX-POOL-INS-PATCH 0.1498 ± 0.0561 0.2932 ± 0.1501 0.5536 ± 0.1972 MEAN-POOL-INS 0.3097 ± 0.0367 0.5338 ± 0.0678 0.7841 ± 0.0780 0.581 ± 0.022 0.589 ± 0.021 63.0 M 525.8 K
3.3 Results
3.3.1 CAMELYON16
The results for this dataset are presented in Tables 1 and 2. In Table 1, the ACMIL/2 model demonstrates moderate classification performance but high reliability. Conversely, DTFD achieves strong classification performance with lower reliability. Interestingly, the MAX-POOL model, while computationally efficient and yielding reasonable classification results, exhibits the lowest reliability. This finding reinforces our argument that model selection should not be based solely on classification metrics. To illustrate this point, we present a tumor slide that was correctly classified by MAX-POOL, along with its heatmap in Figure LABEL:fig:cam_max. The heatmap reveals that the model concentrates on normal patches of lymph node tissue rather than metastatic tumor patches, indicating the unreliability of its predictions.
The results in Table 2 show that additive models do not exhibit significant improvements in classification performance, but they do show slight enhancements in reliability metrics. Notably, the simple MEAN-POOL-INS model achieves the highest reliability values, although having the lowest classification performance. Figure LABEL:fig:cam_mean illustrates a slide and its corresponding heatmap, highlighting the model’s focus on tumor areas.
[…]
fig:cam_mean Refer to caption        Refer to caption
Figure 3: The test-40 slide with ground truth annotation (green) along with its heatmap generated by MEAN-POOL-INS.
Table 3: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for CATCH.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.3015 ± 0.0185 0.5346 ± 0.0090 0.9103 ± 0.0071 0.987 ± 0.002 0.902 ± 0.015 94.4 M 791.3 K CLAM 0.3234 ± 0.0063 0.5383 ± 0.0036 0.9139 ± 0.0021 0.987 ± 0.002 0.881 ± 0.014 94.4 M 798.4 K DTFD 0.2007 ± 0.0599 0.5616 ± 0.0077 0.9141 ± 0.0054 0.995 ± 0.001 0.930 ± 0.011 126.1 M 1841.9 K ACMIL/4 0.2812 ± 0.0108 0.5546 ± 0.0083 0.9180 ± 0.0055 0.995 ± 0.001 0.926 ± 0.010 94.5 M 806.4 K MADMIL/2 0.2983 ± 0.0126 0.5500 ± 0.0042 0.9178 ± 0.0021 0.985 ± 0.001 0.904 ± 0.024 78.7 M 660.2 K MAX-POOL 0.0033 ± 0.0002 -0.0494 ± 0.0032 0.6540 ± 0.0041 0.998 ± 0.000 0.943 ± 0.015 62.9 M 528.3 K MEAN-POOL - - - 0.983 ± 0.001 0.809 ± 0.016 62.9 M 528.3 K
3.3.2 CATCH
Tables 3 and 4 present the results for CATCH dataset. Notably, the MAX-POOL model achieves the highest classification metrics but records the lowest reliability scores. Selecting the MAX-POOL model without considering its reliability could lead to unpredictable outcomes. As shown in Figure LABEL:fig:cat_max, the model primarily focuses on non-tumorous regions, further highlighting its lack of reliability.
Additionally, different models excel in various reliability metrics. Among them, the ACMIL/4 models (both plain and additive) strike a good balance between classification and reliability performance, although they are computationally demanding. In Figure LABEL:fig:cat_acmil, the heatmap produced by the plain ACMIL/4 model demonstrates that the model accurately focuses on regions of interest as annotated by pathologists.
Finally, the MEAN-POOL-INS model, while being the most computationally efficient, exhibits the lowest F1 score but achieves moderate reliability metrics
[…]
fig:cat_max Refer to caption        Refer to caption
Figure 4: A slide from CATCH with ground truth annotation (green) along with its heatmap generated by MAX-POOL.
Table 4: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CATCH using additive models.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.2948 ± 0.0197 0.5266 ± 0.0030 0.9044 ± 0.0049 0.983 ± 0.003 0.896 ± 0.017 94.8 M 791.3 K ABMIL-ADD-PATCH 0.0589 ± 0.0068 0.3465 ± 0.0153 0.8253 ± 0.0073 CLAM-ADD-ATT 0.3077 ± 0.0218 0.5278 ± 0.0080 0.9014 ± 0.0050 0.983 ± 0.003 0.890 ± 0.016 94.8 M 798.5 K CLAM-ADD-PATCH 0.0672 ± 0.0023 0.3467 ± 0.0139 0.8262 ± 0.0056 DTFD-PATCH 0.0920 ± 0.0047 0.4187 ± 0.0085 0.8678 ± 0.0041 0.995 ± 0.001 0.930 ± 0.011 126.1 M 1841.9 K ACMIL/4-ADD-ATT 0.2280 ± 0.0796 0.5455 ± 0.0115 0.9086 ± 0.0066 0.993 ± 0.001 0.916 ± 0.014 94.9 M 806.4 K ACMIL/4-ADD-PATCH 0.1726 ± 0.0121 0.4438 ± 0.0169 0.8783 ± 0.0063 MAX-POOL-INS 0.0004 ± 0.0001 0.0043 ± 0.0068 0.6645 ± 0.0037 0.987 ± 0.005 0.885 ± 0.006 63.3 M 528.4 K MAX-POOL-INS-PATCH 0.2926 ± 0.0116 0.4720 ± 0.0581 0.8764 ± 0.0298 MEAN-POOL-INS 0.3194 ± 0.0034 0.4521 ± 0.0077 0.8525 ± 0.0039 0.983 ± 0.001 0.809 ± 0.016 63.3 M 528.4 K
[…]
fig:cat_acmil Refer to caption        Refer to caption
Figure 5: A slide from CATCH with ground truth annotation (green) along with its heatmap generated by ACMIL/4.
3.3.3 TCGA BRCA
The MADMIL/3 models exhibit the highest reliability performance, a high AUC classification metric, and low computational costs for this dataset, as shown in Table 5. Figure LABEL:fig:brca_madmil visually supports these numerical results. On the other hand, the DTFD model achieves the highest F1 score but demonstrates lower reliability, AUC, and computational efficiency. Additionally, we present a slide’s heatmap generated by the MEAN-POOL-INS model in Figure LABEL:fig:brca_mean-pool-ins, which aligns with the model’s moderate reliability, high classification AUC, and low computational cost reported in Table 6.
Table 5: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for TCGA BRCA.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.1858 ± 0.0384 0.4358 ± 0.0360 0.8206 ± 0.0163 0.953 ± 0.008 0.853 ± 0.022 94.4 M 788.7 K CLAM 0.1779 ± 0.0138 0.4844 ± 0.0129 0.8388 ± 0.0088 0.951 ± 0.007 0.835 ± 0.011 94.4 M 790.7 K DTFD 0.1524 ± 0.0313 0.4748 ± 0.0104 0.8380 ± 0.0081 0.955 ± 0.006 0.869 ± 0.021 125.8 M 1052.6 K ACMIL/3 0.2211 ± 0.0119 0.4908 ± 0.0184 0.8498 ± 0.0113 0.961 ± 0.006 0.852 ± 0.018 94.5 M 792.3 K MADMIL/3 0.2238 ± 0.0155 0.4928 ± 0.0254 0.8530 ± 0.0164 0.966 ± 0.008 0.847 ± 0.040 73.5 M 614.8 K MAX-POOL 0.0038 ± 0.0003 -0.0639 ± 0.0045 0.6109 ± 0.0045 0.932 ± 0.003 0.796 ± 0.023 62.9 M 525.8 K MEAN-POOL - - - 0.977 ± 0.006 0.855 ± 0.017 62.9 M 525.8 K
[…]
fig:brca_madmil Refer to caption        Refer to caption
Figure 6: A slide from TCGA BRCA with ground truth annotation (green) along with corresponding heatmap generated by MADMIL/3.
Table 6: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for TCGA BRCA using additive models.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.1907 ± 0.0256 0.4091 ± 0.1029 0.8116 ± 0.0418 0.910 ± 0.036 0.798 ± 0.035 94.5 M 788.7 K ABMIL-ADD-PATCH 0.0746 ± 0.0287 0.3265 ± 0.0264 0.8034 ± 0.0149 CLAM-ADD-ATT 0.1423 ± 0.0379 0.4770 ± 0.0242 0.8385 ± 0.0179 0.917 ± 0.011 0.786 ± 0.063 94.5 M 790.7 K CLAM-ADD-PATCH 0.0621 ± 0.0187 0.2863 ± 0.0377 0.7889 ± 0.0021 DTFD-PATCH 0.0667 ± 0.0085 0.2949 ± 0.0162 0.7942 ± 0.0058 0.955 ± 0.006 0.869 ± 0.021 125.8 M 1052.6 K ACMIL/3-ADD-ATT 0.1248 ± 0.0122 0.2254 ± 0.2965 0.7324 ± 0.1197 0.937 ± 0.010 0.843 ± 0.020 94.6 M 792.3 K ACMIL/3-ADD-PATCH 0.0587 ± 0.0056 0.1125 ± 0.0750 0.7026 ± 0.0391 MAX-POOL-INS 0.0002 ± 0.0001 0.0104 ± 0.0070 0.6132 ± 0.0043 0.894 ± 0.040 0.725 ± 0.057 63.0 M 525.8 K MAX-POOL-INS-PATCH 0.2364 ± 0.0427 0.4495 ± 0.0403 0.8392 ± 0.0165 MEAN-POOL-INS 0.2672 ± 0.0027 0.3037 ± 0.0107 0.7873 ± 0.0054 0.976 ± 0.006 0.850 ± 0.018 63.0 M 525.8 K
[…]
fig:brca_mean-pool-ins Refer to caption        Refer to caption
Figure 7: A slide from TCGA BRCA with ground truth annotation (green) along with corresponding heatmap generated by MEAN-POOL-INS.
Table 7: The overall mean of the average reliability, classification, and computation metrics across CAMELYON16, CATCH, and TCGA BRCA.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.2013 0.4195 0.7719 0.929 0.866 94.4 M 789.5 K CLAM 0.2063 0.4384 0.7805 0.932 0.855 94.4 M 793.2 K DTFD 0.1602 0.4434 0.7859 0.947 0.882 125.9 M 1315.7 K ACMIL 0.2211 0.4602 0.7844 0.926 0.862 94.4 M 796.5 K MADMIL 0.2183 0.4495 0.7929 0.937 0.866 76.9 M 644.2 K MAX-POOL 0.0029 -0.0377 0.4967 0.909 0.844 62.9 M 526.6 K MEAN-POOL - - - 0.847 0.751 62.9 M 526.6 K
3.4 Discussion
Based on the overall mean of the average results across the three datasets, as presented in Tables 7, 8, and Figure LABEL:fig:bar_plot, several key conclusions can be drawn. First, while the simple MAX-POOL network achieves high classification metrics and low computational cost, it demonstrates poor reliability values, making it unsuitable for real-world applications. In contrast, the MEAN-POOL-INS model shows high reliability, acceptable classification performance, and low computational cost, making it a more favorable choice.
The multihead networks, ACMIL and MADMIL, offer both high reliability and classification performance, but at a higher computational expense, particularly ACMIL. Therefore, they are suitable to have a fair balance among different metrics. Meanwhile, the DTFD model delivers the highest classification performance, though this comes at the cost of increased computational demands and reduced reliability. Therefore, when classification performance is a priority, and computational cost and reliability are secondary concerns, DTFD may be a suitable choice. However, in scenarios where computational efficiency and reliability are more critical than classification performance, the MEAN-POOL-INS model is recommended.
Regarding the additive models, while they do not exhibit a significant improvement in classification metrics, they do show a subtle increase in reliability scores, albeit with slightly higher computational cost. Although additive models tend to be more interpretable, they are not necessarily more reliable.
At the end, we urge researchers to include computational cost and reliability metrics alongside classification results when evaluating and developing new models. Our findings demonstrate that classification metrics alone do not fully reflect the diverse capabilities of models, as evidenced by the results in this study.
3.5 Conclusion
In this paper, we proposed a novel approach to comparing MIL models in terms of reliability using three distinct metrics. Notably, we found that the MEAN-POOL-INS model achieves high reliability despite its simple architecture and low computational cost. Additionally, the multihead models, while slightly more computationally expensive, demonstrate higher classification performance and maintain fair reliability.
We hope that future research will incorporate reliability and computational metrics, leading to the development of more dependable and efficient models for WSI classification.
3.6 Acknowledgments
This work was done as a part of the IMI BigPicture project (IMI945358).
Table 8: The overall mean of the average reliability, classification, and computation metrics for additive models across CAMELYON16, CATCH, and TCGA BRCA datasets.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.2063 0.4178 0.7770 0.918 0.840 94.6 M 789.5 K ABMIL-ADD-PATCH 0.0571 0.3035 0.7045 CLAM-ADD-ATT 0.1884 0.4324 0.7789 0.921 0.834 94.6 M 793.3 K CLAM-ADD-PATCH 0.0539 0.2908 0.6992 DTFD-PATCH 0.0651 0.3343 0.7309 0.947 0.882 125.9 M 1315.7 K ACMIL-ADD-ATT 0.1755 0.3724 0.7537 0.905 0.856 94.6 M 796.5 K ACMIL-ADD-PATCH 0.0858 0.1739 0.6227 MAX-POOL-INS 0.0003 0.0206 0.4921 0.882 0.791 63.1 M 526.6 K MAX-POOL-INS-PATCH 0.2262 0.4049 0.7564 MEAN-POOL-INS 0.2987 0.4298 0.8079 0.846 0.749 63.1 M 526.6 K